# Towards Distillation-Resistant Large Language Models: An Information-Theoretic Perspective

### v1

@@ Appendix A Proofs
$\mathcal{I}(X;Z|Y)\geq\mathcal{I}(X;Z^{\prime}|Y).$
@@
where Equation 10 follows from the fact that the mapping from $Z$ to $Z^{\prime}$ is deterministic, Equation 11 applies the chain rule of mutual information, and Equation 12 follows from the non-negativity of mutual information. ∎
@@
The conditional mutual information $\mathcal{I}(X;Z|Y)$ is a special case of Information Bottleneck objective with $\beta=1$ in Equation 3, i.e.,
@@
According to Theorem 2, $Y\to X\to Z$ forms a Markov chain, implying that
@@ B.2 Knowledge Distillation Configuration
We train the teacher model for five epochs and select the final matrix that minimizes $\mathcal{L}_{\text{grad}}$ on the premise of preserving the teacher’s original performance. The distillation loss coefficient $\alpha$ is set to $0.5$ by default, which assigns equal importance to loss terms $\mathcal{L}_{\text{SFT}}$ and $\mathcal{L}_{\text{KD}}$. Following established practices in prior distillation studies (Wang et al., 2025; Gu et al., 2024; Ko et al., 2024), we search distillation hyperparameters within their recommended ranges to reach better performance under clean distillation while adopting results with more prominent defense effectiveness. Specifically, we vary the number of training epochs in {1, 2, 5}. The learning rate is selected from {1e-6, 5e-6, 1e-5, 5e-5} for the Qwen family and {1e-5, 5e-5, 1e-4} for the Llama family. All models are optimized using the AdamW optimizer with a weight decay of $0.01$ and momentum coefficients $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate schedule consists of a linear warm-up over the first 10% of training steps, followed by cosine annealing to zero.
We adopt greedy decoding for all experiments for stable and reproducible evaluations. The max new token is set to 512 by default, and is 2048 for the MATH-500 benchmark as it requires a substantially longer reasoning process.
Appendix C The Pseudocode of the Proposed Defense
Algorithm 1 Training Defensive Matrices
[…]
0: Trained defensive matrices $\mathbf{A}$ and $\mathbf{B}$;
2:for iteration =1=1 to NN do
3:  Sample a batch of data $(\bm{q},\bm{o})\sim\mathcal{D}$;
5:  Obtain teacher logits $\bm{z}_{\text{T}}\leftarrow f_{\text{T}}(\bm{q})$;
6:  $\bm{z}^{\prime}_{\text{T}}\leftarrow(\mathbf{E+AB})\bm{z}_{\text{T}}$;
9:  Compute the overall training loss $\mathcal{L}_{\mathcal{M}}$ via Eq. (LABEL:eq:final-objective) using $\bm{q},\bm{o},\bm{g},\bm{g}^{\prime}$;
10:  Back-propagate and update $\mathbf{A}$ and $\mathbf{B}$;
11:end for
12:Return $\mathbf{A},\mathbf{B}$
Appendix D Additional Results
Table 4: Distillation performance of surrogate student models Qwen2.5-3B-Instruct and Llama-3.2-3B-Instruct under different datasets and distillation algorithms (%).
@@ Appendix D Additional Results
Table 5: Distillation performance of surrogate Qwen2.5-3B-Instruct model under MATH-500 (%).
@@
This part provides experiments on more scenarios.
Distillation of surrogate students. In the main text, we provide distillation results on the black-box student models. We supplement the defense performance on surrogate student models for matrix learning, namely Qwen2.5-3B-Instruct and Llama-3.2-3B-Instruct for the Qwen and Llama series, respectively. The results are shown in Table 4 and Table 5. As expected, the proposed defense performs effectively in this white-box setting and generally exhibits stronger defensive effectiveness than in the black-box scenario.
As observed, the defense leads to an average accuracy drop of $8.79\%$ across various distillation algorithms and surrogate students on GSM8K, which significantly exceeds the drop of $4.52\%$ observed on black-box students, as the transformation matrix is directly optimized against these surrogate models. A similar trend is observed on MMLU and MATH-500. For example, it leads to an average decrease of $6.55\%$ on MATH-500, exceeding the $5.55\%$ reduction under black-box students.
Ablation on the $\mathcal{L}_{\mathrm{CE}}$. We then analyze the influence of $\mathcal{L}_{\mathrm{CE}}$ on the matrix performance. Specifically, $\mathcal{L}_{\mathrm{CE}}$ encourages teacher logits $Z$ to include more information about the ground-truth label $Y$, which improves accuracy while suppressing additional contextual information. We design a variant that removes $\mathcal{L}_{\mathrm{CE}}$ from the training objective to investigate its influence on teacher performance.
Figure 6 reveals that by canceling $\mathcal{L}_{\mathrm{CE}}$, the teacher’s accuracy drops drastically after equipped with the matrix $\mathcal{M}$, highlighting that the defense would be completely unusable in practice without this term. This confirms the rationality of our theory-induced loss design in preserving the teacher acurracy during matrix optimization.
@@
Ablation on the hyperparameter $\alpha$. In existing KD studies, the loss coefficient $\alpha$ is typically set to 0.5 by default. To comprehensively evaluate our defense efficacy, we consider more general settings with varying values of $\alpha$. We adopt the SOTA KD algorithm ABKD as a representative attack and conduct experiments on Qwen2.5-1.5B using GSM8K.
@@
Figure 7 shows that the proposed defense exhibits strong effectiveness across various $\alpha$ values, validating the generalizability of the designed transformation matrix.
Appendix E Visualization Results
@@
Figure 8: Qualitative comparison between the teacher before and after applying the defensive matrix (Llama, GSM8K).
@@
Figure 9: Qualitative comparison between the teacher before and after applying the defensive matrix (Qwen, MATH-500).
@@
Figure 10: Qualitative comparison between the student before and after applying the defensive matrix (Qwen, GSM8K).
@@
Figure 11: Qualitative comparison between the student before and after applying the defensive matrix (Llama, GSM8K).
@@ B. Kantianism
Figure 12: Qualitative comparison between the student before and after applying the defensive matrix (Qwen, MMLU).
@@
Figure 13: Qualitative comparison between the student before and after applying the defensive matrix (Qwen, MATH-500).
@@
Figure 14: Qualitative comparison between the student before and after applying the defensive matrix (Llama, MMLU).

### v2

@@ Appendix A Proofs
$\mathcal{I}(X;Z|Y)\geq\mathcal{I}(X;Z^{\prime}|Y).$ (13)
@@
where eq.˜14 follows from the fact that the mapping from $Z$ to $Z^{\prime}$ is deterministic, eq.˜15 applies the chain rule of mutual information, and eq.˜16 follows from the non-negativity of mutual information. ∎
@@
The conditional mutual information $\mathcal{I}(X;Z|Y)$ is a special case of Information Bottleneck objective with $\beta=1$ in eq.˜3, i.e.,
@@
According to theorem˜2, $Y\to X\to Z$ forms a Markov chain, implying that
@@ B.2 Knowledge Distillation Configuration
We train the transformation matrix for $5$ epochs and select the final matrix that minimizes $\mathcal{L}_{\text{grad}}$ on the premise of preserving the teacher’s original performance. The distillation loss coefficient $\alpha$ is set to $0.5$ by default, which assigns equal importance to loss terms $\mathcal{L}_{\text{SFT}}$ and $\mathcal{L}_{\text{KD}}$. Following established practices in prior distillation studies Wang et al. (2025); Gu et al. (2024); Ko et al. (2024), we search distillation hyperparameters within their recommended ranges to reach the best performance under clean distillation, and report the corresponding performance after applying our defense. Specifically, the learning rate is selected from {1e-6, 5e-6, 1e-5, 5e-5} for the Qwen family and {1e-5, 5e-5, 1e-4} for the Llama family. All models are optimized using the AdamW optimizer with a weight decay of $0.01$ and momentum coefficients $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate schedule consists of a linear warm-up over the first 10% of training steps, followed by cosine annealing to zero. We adopt greedy decoding for all experiments for stable and reproducible evaluations. The max new token is set to 512 by default, and is 2048 for the MATH-500 benchmark as it requires a substantially longer reasoning process. All the experiments are performed with 2 NVIDIA A800 GPUs.
Appendix C Additional Results
Distillation of surrogate students. In the main text, we present distillation results on black-box student models to evaluate the practical effectiveness of our defense under realistic deployment scenarios. Here, we further supplement the analysis by reporting defense performance on surrogate student models used during matrix learning, namely Qwen2.5-3B-Instruct for the Qwen series and Llama-3.2-3B-Instruct for the Llama series.
The corresponding results are summarized in Table 4 and Table 4. As expected, the proposed defense demonstrates strong effectiveness in this white-box setting and generally exhibits stronger defensive effectiveness than in the black-box scenario.
Table 3: Distillation performance of surrogate students Qwen2.5-3B-Instruct and Llama-3.2-3B-Instruct under different distillation algorithms (%).
@@ Appendix C Additional Results
Table 4: Distillation performance of surrogate Qwen2.5-3B-Instruct under MATH-500 (%).
@@
As observed, the proposed defense results in an average accuracy drop of $8.79\%$ across various distillation algorithms and surrogate student models on GSM8K, which significantly exceeds the $4.52\%$ reduction observed for black-box students. This discrepancy arises because the transformation matrix is explicitly optimized against these surrogate models, leading to stronger and more targeted defensive effects. A similar trend can be consistently observed on MMLU and MATH-500. For instance, on MATH-500, the defense yields an average performance decrease of $6.55\%$, which again surpasses the $5.55\%$ reduction observed under black-box settings. These results further demonstrate the effectiveness and generalizability of our approach across different benchmarks and threat scenarios.
Ablation on the $\mathcal{L}_{\mathrm{CE}}$. We then analyze the influence of $\mathcal{L}_{\mathrm{CE}}$ on the matrix performance. Specifically, $\mathcal{L}_{\mathrm{CE}}$ encourages the teacher logits $Z$ to encode more information about the ground-truth label $Y$, improving predictive accuracy while suppressing redundant contextual information. To further investigate its role, we design a variant that removes $\mathcal{L}_{\mathrm{CE}}$ from the training objective and examine its impact on teacher performance. As shown in Figure 7, eliminating $\mathcal{L}_{\mathrm{CE}}$ causes the teacher’s accuracy to drop drastically after being equipped with the transformation matrix $\mathcal{M}$, rendering the defense practically unusable. This observation highlights the critical importance of $\mathcal{L}_{\mathrm{CE}}$ and further validates the rationality of our theoretically grounded loss design in effectively preserving teacher accuracy during matrix optimization.
Ablation on the hyperparameter $\alpha$. In existing KD studies, the coefficient $\alpha$ is typically set to $0.5$. To more comprehensively evaluate the efficacy of our defense, we further consider more general settings with varying $\alpha$ values. We adopt the SOTA KD algorithm ABKD as a representative attack and conduct experiments on Qwen2.5-1.5B using GSM8K. As in Figure 7, the proposed defense consistently demonstrates strong effectiveness across diverse choices of $\alpha$, validating the robustness and generalizability of the designed transformation matrix under varying distillation configurations.
@@
Appendix D The Pseudocode of the Proposed Defense
Algorithm 1 Training Defensive Matrices
0:Input: $f_{\text{T}}(\cdot)$: the teacher model; $f_{\text{S}}(\cdot)$: the proxy student model with parameter $\theta$; $\mathcal{D}$: the training dataset; $|\mathcal{V}|$: the vocabulary size; $r$: the rank; $\lambda$: the loss coefficient; NN: the number of training iterations;
0:Output: Trained defensive matrices $\mathbf{A}$ and $\mathbf{B}$;
2:for iteration =1=1 to NN do
3:  Sample a batch of data $(\bm{q},\bm{o})\sim\mathcal{D}$;
5:  Obtain teacher logits $\bm{z}_{\text{T}}\leftarrow f_{\text{T}}(\bm{q})$;
6:  $\bm{z}^{\prime}_{\text{T}}\leftarrow(\mathbf{E+AB})\bm{z}_{\text{T}}$;
9:  Compute the overall training loss $\mathcal{L}_{\mathcal{M}}$ via Eq. (11) using $\bm{q},\bm{o},\bm{g},\bm{g}^{\prime}$;
10:  Back-propagate and update $\mathbf{A}$ and $\mathbf{B}$;
11:end for
12:Return $\mathbf{A},\mathbf{B}$
Appendix E Limitations
Despite the encouraging experimental results, this work still faces several limitations. First, due to the high-dimensional and intractability of text distributions, we do not provide a rigorous theoretical proof that optimizing $\mathcal{L}_{grad}$ is equivalent to maximizing the mutual information $\mathcal{I}(X;Z)$. However, the proposed gradient angle maximization strategy tightly align with the underlying mechanism of $\mathcal{I}(X;Z)$ maximization, which has also been experimentally proven to be effective. Future work can investigate this issue and propose more rigorous theoretical foundations. Nevertheless, this overhead remains modest in practice, as the number of parameters introduced by the matrix is relatively small (only about 0.1% of the total model parameters). Consequently, the overall impact on inference efficiency is limited and does not significantly affect deployment.
Appendix F Visualization Results
@@
Figure 8: Qualitative comparison between the teacher before and after applying the defensive matrix with Llama-3.1-8B on GSM8K.
@@
Figure 9: Qualitative comparison between the teacher before and after applying the defensive matrix with Qwen2.5-7B on MATH-500.
@@
Figure 10: Qualitative comparison between the student before and after applying the defensive matrix with Qwen2.5-1.5B on GSM8K.
@@
Figure 11: Qualitative comparison between the student before and after applying the defensive matrix with Llama-3.2-1B on GSM8K.
@@ B. Kantianism
Figure 12: Qualitative comparison between the student before and after applying the defensive matrix with Qwen2.5-1.5B on MMLU.
@@
Figure 13: Qualitative comparison between the student before and after applying the defensive matrix with Qwen2.5-1.5B on MATH-500.
@@
Figure 14: Qualitative comparison between the student before and after applying the defensive matrix with Llama-3.2-1B on MMLU.